SpaceRef

SpaceRef


SOHO Daily Meeting Minutes for Friday 20 June 2003

Status Report From: Goddard Space Flight Center
Posted: Friday, June 20, 2003

ANNOUNCEMENTS

* See note at end for latest HGA information.

* There was no weekly meeting today. See note at end for details on next week.

* A Major Flare Watch has been declared on AR10386. The campaign number is 6961.

* EIT Shutterless is today. No NRT will be available from 17-19:40 UT.

* GITT scans for the EOF and EAF are scheduled for Thursday 26 June at 9:30am.

FOT REPORT

Spacecraft Anomalies: None

Accomplished Activities: HGA test activities

Planned Activities: VIRGO, SWAN, MDI, RSL, New tracking star, mag = 7.5

Upcoming Operations: None

Ground Anomalies: None

SOLAR STATUS

06/19 DOY 170 - 06/20 DOY 171

EIT daily On-Line Solar Status and observations: http://umbra.nascom.nasa.gov/eit/plan/log_form20030620.html

LASCO daily On-Line Solar Status and observations: http://lasco-www.nrl.navy.mil/observations/Todays_obs.html

Web Page for Planning: http://www.bbso.njit.edu/arm

Magnetic Maps and Heliospheric Forecasts: http://www.lmsal.com/forecast/

INSTRUMENT STATUS

CDS: Nominal. JOP153, EIT shutterless support, La Palma study, coronal hole spectral atlas. Weekend: Explosive events study, off-limb spectral study in the north and south, GIS spectral atlas, CME onset study, engineering, GIS QS monitoring.

UVCS: Nominal. Streamer study and synoptics.

LASCO: Nominal. C2/C3 synoptics, JOP153, JOP159.

EIT: Nominal. 195 CME watch and synoptics, JOP153, JOP159, and shutterless campaign.

MDI: Nominal. Full disk magnetograms and dopplergrams, JOP153, EIT shutterless and La Palma support.

TRACE: Nominal. La Palma support, JOP153.

AOB

Yesterday morning we tried several options to move the HGA, without success. We then decided to use both the A and B motor windings simultaneously, in an effort to increase the motor torque. This was successful and we managed to move the antenna by +4.2 deg to the "sweet spot" at -17.9 deg. 4 encoder pulses were received at exactly the right time. An attempt of operating the motor with only the A side winding after the first two pulses in the double-motor mode was unsuccessful again. In this position, we would lose about 19 days of high rate telemetry every three months, starting late next week. This would give us approximately 80% coverage.

Starting mid-week next week we expect to lose the HGA signal for ~20 days. Instruments need to prepare for this outage on Monday and Tuesday.

// end //

More status reports and news releases or top stories.

Please follow SpaceRef on Twitter and Like us on Facebook.

NanoRacks